# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v1

@@ V Conclusion
V-A Proof of Theorem 1
Proof Outline: To prove Theorem 1, we first derive four lemmas, namely Lemmas 1$\sim$4, to establish the relationship between the convergence rate and model error. Then Lemma 5 builds the connection between model error and BER. By combining these relations, Theorem 1 is proved, which presents the relationship between convergence rate and BER.
For each communication round $r=1,...,R$ and local iteration $e=1,...,E$, we have
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{s}\sum_{k\in{S_{r}}}(\mathbf{w}_{r,E}^{k}-(Q(\mathbf{w}_{r}))^{\prime}),$
$\hat{\mathbf{w}}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in{[n]}}(\mathbf{w}_{r,E}^{k}-(Q(\mathbf{w}_{r}))^{\prime}),$
$\bar{\mathbf{w}}_{r,e}=\frac{1}{n}\sum_{k\in{[n]}}\mathbf{w}_{r,e}^{k},$ (14)
where $\mathbf{w}_{r+1}$ represents the updated global model based on $s$ selected clients, $\hat{\mathbf{w}}_{r+1}$ denotes the updated global model based on all $n$ clients, and $\bar{\mathbf{w}}_{r,e}$ denotes the average model of $n$ local models. $(Q(\mathbf{w}_{r}))^{\prime}$ represents that the downlink model $\mathbf{w}_{r}$ will be impacted by the quantization operation $Q(\cdot)$ and channel noise $(\cdot)^{\prime}$ before being received by the clients. To simplify the representation, we will use $Q^{\prime}(\cdot)$ to replace $(Q(\cdot))^{\prime}$ in subsequent derivations.
Lemma 1. Assuming that Assumptions 1 and 4 hold, for any communication round $r=1,...,R$, we have
$\mathbb{E}f(\mathbf{w}_{r+1})$ $\leq\mathbb{E}f(\bar{\mathbf{w}}_{r,E})+\frac{L}{2}\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}$
$+\frac{L}{2}\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\mathbf{w}_{r+1}\|^{2}.$ (15)
This lemma provides a bound on the training loss of the updated global model. This bound can be derived from the $L$-smooth property of the loss function on the server side. The proof is given in Appendix-B.
In the next three lemmas, we derive the upper bounds for the three terms on the RHS of (15).
Lemma 2. Assuming that Assumptions 1 - 3 hold and considering the sequence of updates with step size $\eta$, the following inequality holds for the case of noisy downlink transmission:
$\mathbb{E}f(\bar{\mathbf{w}}_{r,E})\leq\frac{L}{2}\mathbb{E}\|Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$+\mathbb{E}f(\mathbf{w}_{r})-\frac{\eta}{2n}(1-L\eta)\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{E\eta^{2}\sigma_{G}^{2}}{2}.$ (16)
The proof is given in Appendix-B. This lemma captures the process of noisy broadcasting and local updates. As shown in the RHS of Eq.(16), the first two terms arise from the noisy broadcasting, while the subsequent terms come from the local training process.
Lemma 3. For $\hat{\mathbf{w}}_{r+1}$ and $\bar{\mathbf{w}}_{r,E}$ defined in (14), we have
$\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}=\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}.$ (17)
This lemma indicates that the expectation of the global model error on the server side is the same as the downlink model error on client side. The proof can be found in Appendix-B.
Lemma 4. For $\mathbf{w}_{r+1}$ and $\hat{\mathbf{w}}_{r+1}$ defined in (14), we have
$\mathbb{E}\|\mathbf{w}_{r+1}-$ $\hat{\mathbf{w}}_{r+1}\|^{2}\leq\frac{4}{s(n-1)}(1-\frac{s}{n})\cdot$
$\left[\eta^{2}\sigma_{L}^{2}En+\eta^{2}E\sum_{e=1}^{E}\sum_{k\in[n]}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\right].$ (18)
This lemma captures the effect of partial participation of clients. The proof is provided in Appendix-B.
By substituting Lemmas 2-4 into Lemma 1, we derive the following recursive inequality for the updated model at the server:
$\mathbb{E}f(\mathbf{w}_{r+1})\leq L\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
@@ V-A Proof of Theorem 1
$-\frac{\eta}{2n}\left\{1-L\eta-\frac{4LE(n-s)}{s(n-1)}\right\}\cdot\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\mathbb{E}f(\mathbf{w}_{r})+\frac{\eta E\sigma_{G}^{2}}{2}+L\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (19)
When $\eta$ satisfies the condition
$1-L\eta-\frac{4LE(n-s)}{s(n-1)}\geq 0,$ (20)
we have
$\mathbb{E}f(\mathbf{w}_{r+1})\leq L\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$+\mathbb{E}f(\mathbf{w}_{r})+\frac{E\eta^{2}\sigma_{G}^{2}}{2}+L\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (21)
Summing (21) over $r=1,...,R$ and rearranging the terms, we obtain the following expression:
$\frac{\eta}{2}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq L\sum_{r=1}^{R}\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$
$+f(\mathbf{w}_{0})-f^{*}+\frac{RE\eta^{2}\sigma_{G}^{2}}{2}+RL\eta^{2}\sigma_{L}^{2}E(\frac{1}{2n}+\frac{2(n-s)}{s(n-1)}).$ (22)
Multiplying $\frac{2}{RE\eta}$ on both sides of (22), we get
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2L}{RE\eta}\sum_{r=1}^{R}\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$
$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (23)
Since a high quantization bit-length is assumed to be used in this work with negligible quantization error, we have
$\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2}$ $=\mathbb{E}\|Q^{\prime}(\mathbf{w}_{r})-Q(\mathbf{w}_{r})+Q(\mathbf{w}_{r})-\mathbf{w}_{r}\|^{2}$
$=\mathbb{E}\|(\mathbf{w}_{r})^{\prime}-\mathbf{w}_{r}\|^{2}.$ (24)
Combining (23) and (24), we obtain
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2L}{RE\eta}\sum_{r=1}^{R}\mathbb{E}\|(\mathbf{w}_{r})^{\prime}-\mathbf{w}_{r}\|^{2}$
$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (25)
Lemma 5. For the perfect global model $\mathbf{w}$ and the received noisy model $(\mathbf{w})^{\prime}$, the model error can be bounded by
$\mathbb{E}||(\mathbf{w})^{\prime}-\mathbf{w}||^{2}\leq\frac{4d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\cdot M^{2},$ (26)
where $M$ denotes the maximum absolute value among all weight elements, i.e. $,M=max(|w_{1}|,|w_{2}|,$$,|w_{d}|)$. Lemma 5 gives the relationship between model variance and BER. The proof can be found in Appendix-B.
By substituting Eq.(26) into Eq.(25), we establish the relationship between the FL convergence rate and BER:
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{8LdM^{2}}{3RE\eta}\sum_{r=1}^{R}\frac{b_{r}(4^{N}-1)}{(2^{N}-1)^{2}}$
$+\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}+\eta\sigma_{G}^{2}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)}).$ (27)
When the quantization bit-length is high, we have $\frac{4^{N}-1}{(2^{N}-1)^{2}}\approx 1$, and Eq.(27) can be simplified as follows:
$\frac{1}{RE}\sum_{r=1}^{R}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}\leq\frac{2(f(\mathbf{w}_{0})-f^{*})}{RE\eta}$
$+\frac{8LdM^{2}}{3RE\eta}\sum_{r=1}^{R}b_{r}+L\eta\sigma_{L}^{2}(\frac{1}{n}+\frac{4(n-s)}{s(n-1)})+\eta\sigma_{G}^{2},$ (28)
which completes the proof of Theorem 1.
V-B Proof of Lemmas
1) Proof of Lemma 1
Quantization is unbiased according to Assumption 3, i.e., $\mathbb{E}_{Q}{(\mathbf{w})}=\mathbf{w}$. It is proved in Eq.(44) that the impact of the BER is also unbiased. So we have $\mathbb{E}_{Q^{\prime}}{Q^{\prime}(\mathbf{w})}=\mathbf{w}$. Based on the above unbiased properties and the notations defined in Eq.(14), we have
$\mathbb{E}_{Q^{\prime}}\hat{\mathbf{w}}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-\mathbb{E}_{Q^{\prime}}Q^{\prime}(\mathbf{w}_{r}))$
$=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})$
$=\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}$
$=\bar{\mathbf{w}}_{r,E}.$ (29)
We can also get
$\mathbb{E}_{S_{r}}\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-Q^{\prime}(\mathbf{w}_{r}))=\hat{\mathbf{w}}_{r+1},$ (30)
whose proof is similar to Section 7.1 in [35].
Since the loss function $f$ is $L$-smooth, for any variables $\mathbf{x,y}$, we have
@@ V-B Proof of Lemmas
$f(\mathbf{x})\leq f(\mathbf{y})+\left\langle\nabla f(\mathbf{y}),\mathbf{x}-\mathbf{y}\right\rangle+\frac{L}{2}||\mathbf{x}-\mathbf{y}||^{2}.$ (31)
Taking $\mathbf{x}=\hat{\mathbf{w}}_{r+1}$ and $\mathbf{y}=\mathbf{\bar{w}}_{r,E}$ in (31), we can get
@@
$f(\hat{\mathbf{w}}_{r+1})\leq$ $f(\mathbf{\bar{w}}_{r,E})+\left\langle\nabla f(\mathbf{\bar{w}}_{r,E}),\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}\right\rangle$
$+\frac{L}{2}||\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}.$ (32)
We take the expectation on both sides of (32) and, since $\mathbb{E}_{Q^{\prime}}\hat{\mathbf{w}}_{r+1}=\bar{\mathbf{w}}_{r,E}$ (see (29)), we have
$\mathbb{E}f(\hat{\mathbf{w}}_{r+1})\leq\mathbb{E}f(\mathbf{\bar{w}}_{r,E})+\frac{L}{2}\mathbb{E}||\hat{\mathbf{w}}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}.$ (33)
Taking $\mathbf{x}=\mathbf{w}_{r+1}$ and $\mathbf{y}=\hat{\mathbf{w}}_{r+1}$ in (31), and following the same steps as in (32) and (33), we have
$\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\hat{\mathbf{w}}_{r+1})+\frac{L}{2}\mathbb{E}||\hat{\mathbf{w}}_{r+1}-\mathbf{w}_{r+1}||^{2}.$ (34)
By combining (33) and (34), we can get Lemma 1.
2) Proof of Lemma 2
Based on the proof in Section 8.3 of [35], and considering the non-iid case as in Assumption 3, if the downlink communication is perfect, we can obtain the upper bound of FL convergence as
$\mathbb{E}$ $f(\bar{\mathbf{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$-\eta(\frac{1}{2n}-\frac{L\eta}{2n}-\frac{L^{2}\eta^{2}E(E-1)}{n})\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.$ (35)
For noisy downlink communication, the received model will have quantization and BER noise. As a result, (35) can be rewritten as
$\mathbb{E}$ $f(\bar{\mathbf{w}}_{r,E})\leq\mathbb{E}f(Q^{\prime}(\mathbf{w}_{r}))-\frac{\eta}{2}\sum_{e=1}^{E}\mathbb{E}\|\nabla f(\bar{\mathbf{w}}_{r,e})\|^{2}$
$-\eta(\frac{1}{2n}-\frac{L\eta}{2n}-\frac{L^{2}\eta^{2}E(E-1)}{n})\sum_{e=1}^{E}\sum_{k\in[n]}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}$
$+\frac{LE\eta^{2}\sigma_{L}^{2}}{2n}+\frac{E\eta^{2}\sigma_{G}^{2}}{2}.$ (36)
Since $f$ is $L$-smooth, according to (31), we have
@@
$f(Q^{\prime}(\mathbf{w}_{r}))\leq$ $f(\mathbf{w}_{r})+\left\langle\nabla f(\mathbf{w}_{r}),Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}\right\rangle$
$+\frac{L}{2}||Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}||^{2}.$ (37)
Since the quantization and BER are both unbiased, i.e., $\mathbb{E}_{Q^{\prime}}{Q^{\prime}(\mathbf{w})}=\mathbb{E}_{Q}{Q(\mathbf{w})}=\mathbf{w}$, we take the expectation on both sides of (37) to obtain
$\mathbb{E}f(Q^{\prime}(\mathbf{w}_{r}))\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}||Q^{\prime}(\mathbf{w}_{r})-\mathbf{w}_{r}||^{2}.$ (38)
Combining (36) and (38), we prove Lemma 2.
3) Proof of Lemma 3
According to the notations defined in Eq.(14), we can get the expanded representation of $\hat{\mathbf{w}}_{r+1}$ and $\bar{\mathbf{w}}_{r,E}$ and derive
$\mathbb{E}\|\hat{\mathbf{w}}_{r+1}-\bar{\mathbf{w}}_{r,E}\|^{2}$
$=\mathbb{E}\|\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}(\mathbf{w}_{r,E}^{k}-Q^{\prime}(\mathbf{w}_{r}))-\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}\|^{2}$
$=\mathbb{E}\|\mathbf{w}_{r}+\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}-\frac{1}{n}\sum_{k\in[n]}Q^{\prime}(\mathbf{w}_{r})-\frac{1}{n}\sum_{k\in[n]}\mathbf{w}_{r,E}^{k}\|^{2}$
$=\mathbb{E}\|\mathbf{w}_{r}-Q^{\prime}(\mathbf{w}_{r})\|^{2},$ (39)
which concludes the proof of Lemma 3.
4) Proof of Lemma 4
Similar to Section 8.5 in [35] and considering the noise in the downlink global model, we have
$\mathbb{E}\|\mathbf{w}_{r+1}-\hat{\mathbf{w}}_{r+1}\|^{2}$ $\leq\frac{4}{s(n-1)}(1-\frac{s}{n})\cdot$
$\sum_{k\in[n]}\mathbb{E}\|\mathbf{w}^{k}_{r,E}-Q^{\prime}(\mathbf{w}_{r})\|^{2}.$ (40)
Based on $E$ steps of local training with learning rate $\eta$, we have

### v2

@@ V Conclusion
V-A Proof of Lemma 1
We express a decoded model $\tilde{\mathbf{w}}$ in terms of $\mathbf{w}$ as follows:
$\tilde{\mathbf{w}}=\mathbf{w}+\bm{\delta}\left(\mathbf{w}\right),$ (18)
where $\bm{\delta}\left(\mathbf{w}\right)$ is a $d$-dimensional function of $\mathbf{w}$ denoting the model distortion caused by bit errors. Under Assumption 4, each entry of $\bm{\delta}\left(\mathbf{w}\right)$ depends only on $w_{d}$, i.e., the $d$-th dimension of $\mathbf{w}$. Denote $\delta\left(w_{d}\right)$ as the $d$-th dimension of $\bm{\delta}\left(\mathbf{w}\right)$. Therefore, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (19)
Under Assumptions 4 and 5, the 1-bit error may occur at different bits, leading to different levels of model distortion. Let $w_{d}^{(i)}\in\{0,1\}$ be the $i$-th digit ($i=0$ and $N-1$ respectively denote the least and most significant bits) in the $N$-bit representation of model parameter $w_{d}$. Hence,
$\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]=w_{d}+b\left(1-b\right)^{N-1}\Delta\sum_{i=0}^{N-1}\left(1-2w_{d}^{(i)}\right)2^{i},$ (20)
The expected bias of $w_{d}$ is bounded as follows:
$\begin{split}|\mathbb{E}\left[\tilde{w}_{d}|w_{d}\right]-w_{d}|&\leq b\left(1-b\right)^{N-1}\left(2^{N}-1\right)\Delta\\ &=b\left(1-b\right)^{N-1}M.\end{split}$ (21)
The mean square error of $w_{d}$ is given as follows:
$\begin{split}\mathbb{E}\left[||\tilde{w}_{d}-w_{d}||^{2}|w_{d}\right]&=b\left(1-b\right)^{N-1}\sum_{i=0}^{N-1}\left(\Delta 2^{i}\right)^{2}\\ &=b\left(1-b\right)^{N-1}\frac{4^{N}-1}{3(2^{N}-1)^{2}}M^{2}.\end{split}$ (22)
Therefore,
@@ V-A Proof of Lemma 1
$\begin{split}\mathbb{E}\left[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}\right]&=\sum_{j=1}^{d}\mathbb{E}\left[||\tilde{w}_{j}-w_{j}||^{2}|\mathbf{w}\right]\\ &=\frac{d(4^{N}-1)}{3(2^{N}-1)^{2}}\cdot b\left(1-b\right)^{N}M^{2}.\end{split}$ (23)
V-B Proof of Lemma 2
We define
$\mathbf{w}_{r+1}$ $=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{\tilde{w}}_{r}),$ (24)
$\mathbf{\bar{w}}_{r,e}$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbf{w}_{r,e}^{k},$ (25)
for each communication round $r=0,...,R-1$ and each local iteration $e=0,\cdots,E-1$, where $\mathbf{w}_{r+1}$ denotes the updated global model and $\mathbf{\bar{w}}_{r,e}$ is the average of client models after the $e$-th local iteration.
To prove the Lemma 2, we first present Lemmas 26$\sim$5.
Lemma 3.
For $\mathbf{\bar{w}}_{r,E}$ and $\mathbf{w}_{r+1}$ defined in (24) and (25), we have
$\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}=\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r},\ r=0,\cdots,R-1.$ (26)
Proof.
The proof can be obtained by substituting the definition of $\mathbf{\bar{w}}_{r,E}$ to the LHS of (26). ∎
Lemma 26 translates the difference between $\mathbf{w}_{r+1}$ and $\mathbf{\bar{w}}_{r,E}$ to the model distortion incurred by downlink bit errors.
Lemma 4.
With Assumption 1, for $r=0,...,R-1$,
@@ V-B Proof of Lemma 2
$\begin{split}\mathbb{E}f(\mathbf{w}_{r+1})&\leq\mathbb{E}f(\mathbf{\bar{w}}_{r,E})+\frac{L}{2}\mathbb{E}||\mathbf{w}_{r+1}-\mathbf{\bar{w}}_{r,E}||^{2}\\ &-\mathbb{E}\left\langle\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle.\end{split}$ (27)
Proof.
The proof can be obtained using the $L$-smoothness of $f(\cdot)$, which is omitted for brevity. ∎
Lemma 27 provides an upper bound on the expected value of the global loss function at $\mathbf{w}_{r+1}$.
Lemma 5.
With Assumptions 1 - 3, for $r=0,\cdots,R-1$,
@@
$\begin{split}&\mathbb{E}f(\mathbf{\bar{w}}_{r,E})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L}{2}\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}\\ &+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (28)
Proof.
The proof can be obtained following similar lines for Lemma 7 in [20] by further considering the non-IID data distribution among clients (i.e., Assumption 3), which is omitted for brevity. ∎
Lemma 28 provides an upper bound on the expected value of the global loss function at $\bar{\mathbf{w}}_{r,E}$ in terms of the expected model distortion caused by downlink bit errors, i.e., $\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}$. By applying Lemmas 26 and 28 to Lemma 27, we obtain the following inequality:
@@
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+L\mathbb{E}\|\mathbf{\tilde{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &+\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &-\frac{\eta}{2n}\left[1-L\eta-2L^{2}\eta^{2}E(E-1)\right]\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}\\ &+\frac{LE\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (29)
By leveraging $2\left\langle\mathbf{a},\mathbf{b}\right\rangle=\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}-\|\mathbf{a}-\mathbf{b}\|^{2}\leq\|\mathbf{a}\|^{2}+\|\mathbf{b}\|^{2}$ holds for any $\mathbf{a}$, $\mathbf{b}$, an upper bound for the third term on the RHS of (29) is derived as follows:
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{\mathbb{E}\|\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E})\|^{2}}{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{a})}{\leq}\frac{L^{2}}{2}\mathbb{E}\|\mathbf{\bar{w}}_{r,E}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\mathbf{w}_{r})\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &=\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{1}{n}\sum_{k=1}^{n}(\mathbf{w}_{r,E}^{k}-\tilde{\mathbf{w}}_{r})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{=}\frac{L^{2}}{2}\mathbb{E}\bigg{\|}\frac{\eta}{n}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\nabla f_{k}(\mathbf{w}_{e}^{k};\xi_{r,e}^{k})+\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\bigg{\|}^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\end{split}$ (30)
where (a) is due to the $L$-smoothness of $f(\cdot)$ and (b) is due to the local updating rules. Based on the Assumptions 2 and 3, we have
$\begin{split}\mathbb{E}\|\nabla&f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}=\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\\ &-\nabla f_{k}(\mathbf{w})+\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})+\nabla f(\mathbf{w})\|^{2}\\ &\leq\sigma_{L}^{2}+\sigma_{G}^{2}+\mathbb{E}\|\nabla f(\mathbf{w})\|^{2}.\end{split}$ (31)
By combing (31) and (30), the bound in (30) can be further updated as
$\begin{split}&\mathbb{E}\left\langle\nabla f(\mathbf{w}_{r})-\nabla f(\mathbf{\bar{w}}_{r,E}),\bm{\delta}(\mathbf{w}_{r})\right\rangle\\ &\leq\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f_{k}(\mathbf{w}_{r,e}^{k};\xi_{r,e}^{k})\|^{2}+\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}\\ &\ \ \ +\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &\overset{(\text{b})}{\leq}\frac{L^{2}\eta^{2}}{2n^{2}}\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}+\frac{L^{2}\eta^{2}E(\sigma_{G}^{2}+\sigma_{L}^{2})}{2n}\\ &\ \ \ +\frac{L^{2}}{2}\mathbb{E}\|\tilde{\mathbf{w}}_{r}-\mathbf{w}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2},\end{split}$ (32)
Substituting (32) in (29), we have
$\begin{split}&\mathbb{E}f(\mathbf{w}_{r+1})\leq\mathbb{E}f(\mathbf{w}_{r})+\frac{L(L+2)}{2}\mathbb{E}\|\mathbf{w}_{r}-\mathbf{\tilde{w}}_{r}\|^{2}+\frac{\mathbb{E}\|\bm{\delta}(\mathbf{w}_{r})\|^{2}}{2}\\ &-\frac{\eta}{2n}\left[1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\right]\\ &\cdot\sum_{k=1}^{n}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{w}_{r,e}^{k})\|^{2}-\frac{\eta}{2}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ &+\frac{L(L+1)E\eta^{2}(\sigma_{L}^{2}+\sigma_{G}^{2})}{2n}+\frac{L^{2}\eta^{3}\sigma_{L}^{2}(n+1)E(E-1)}{2n}.\end{split}$ (33)
For any $\eta$ that satisfies (8), we have